OpenStudy (anonymous):

What are the odds of a 1/5 roll not displaying a 1 after 30 rolls?

OpenStudy (paxpolaris):

like a 5 sided dice huh?

OpenStudy (anonymous):

sure

OpenStudy (paxpolaris):

the odds of not getting a 1 on the first try: \[\Large \frac 45\] the odds of not getting a 1 on the 2nd try: \[\Large \frac 45\times\frac 45\]

OpenStudy (anonymous):

\[\left(\frac{4}{5}\right)^{30}=\frac{1152921504606846976}{931322574615478515625}=0.001238 \]

OpenStudy (anonymous):

so its (4/5)^n (let n be the number of rolls?)

OpenStudy (anonymous):

ok forgive me, but how do i calculate out ^n ... i mean physically how is that calculated

OpenStudy (anonymous):

(4^30)/ (5^30) ?

OpenStudy (paxpolaris):

sure

OpenStudy (shane_b):

Yes

OpenStudy (anonymous):

Ok ty very much for the help guys

OpenStudy (anonymous):

@robtobey's answer was calculated using the Mathematica program which provides unlimited precision integer arithmetic.

OpenStudy (paxpolaris):

@PaxPolaris thanks @robtobey for this information

OpenStudy (anonymous):

You are welcome.